# MTP: Advancing Remote Sensing Foundation Model via Multi-Task Pretraining

### v1

@@ II-C Multi-Task Pretraining for RS Foundation Model
Applying multi-task learning to enhance the RS foundation model is an intuitive idea. Li et al. [64] introduced multi-task SSL representation learning, combining image inpainting, transform prediction, and contrast learning to boost semantic segmentation performance in RS images. However, it was limited to finetuning a pretrained model solely on semantic segmentation tasks, constrained by model size and pretraining dataset capacity. The aspiration to consolidate multiple tasks into a single model has been a longstanding pursuit [15, 17, 65, 66, 67, 68, 69, 70, 71, 58, 72, 73, 42, 74], aligning with the original goals of the foundation model exploration. Bastani et al. [59] devised a multi-task model by integrating Swin-Base [57] with seven heads from existing networks (e.g., Faster-RCNN [75] and UNet [76]), facilitating training on the multi-task annotated Satlas dataset. However, their approach lacked incorporation of typical RS rotated object tasks, focusing solely on transferring the model to RS classification datasets. Inspired by these pioneering efforts, this paper presents multi-task pretraining of RS foundation models with over 300M parameters, encompassing semantic segmentation, instance segmentation, and rotated object detection tasks using the SAMRS dataset. After pretraining, the backbone network is further finetuned on various RS downstream tasks.
@@ III-B Backbone Network
In this research, we adopt RVSA [30] and InternImage [81] as the representative vision transformer-based and CNN-based foundation models.
@@ III-B1 RVSA
This model is specially designed for RS images. Considering the various orientations of RS objects caused by the bird’s-eye view, this model extends the varied-size window attention in [82] by additionally introducing a learnable angle factor, offering windows that can freely zoom, translate, and rotate. RVSA is used to replace the original full attention in original vision transformers. To achieve a trade-off between accuracy and efficiency, following [83], only the full attention in 1/4 depth layer is preserved. In the original paper [30], RVSA is separately used on ViT [84] and ViTAE [71], whereas ViTAE is a CNN-Transformer hybrid model. In this paper, we employ the ViT-based version to investigate the impact of MTP on a plain vision transformer. In addition, the RVSA model in the original paper is limited to the base version of vision transformers, i.e., ViT-B + RVSA. To pretrain larger models, we further apply RVSA to ViT-Large, obtaining ViT-L + RVSA. Their detailed configurations are presented in Table I.
@@ III-B2 InternImage
III-C Implementation of Multi-Task Pretraining
@@
IV Experiments
In this section, we thoroughly evaluate MTP’s performance by finetuning pretrained models across four classical RS tasks: scene classification, object detection, semantic segmentation, and change detection. We also investigate the characteristics of MTP-based RS foundation models, examining the relationships between adopted datasets, hyperparameters, and finetuning performances, measuring accuracy variations with reduced training samples, and visualizing the predicted results.
@@ IV-A4 Finetuning Results and Analyses
Table IV shows the finetuning results. It can be seen that MTP can improve existing foundation models on scene classification tasks, especially for the RVSA series. It helps the model achieve state-of-the-art performances compared to other pretraining models that have comparable sizes. With the help of MTP, on the RESISC-45 dataset, InterImage-XL surpasses Swin-L-based SkySense [42], which is pretrained on a tremendously large dataset that has more than 20 million multimodal RS image triplets involving RGB high-resolution images and multi-temporal multispectral and SAR sequences. MTP boosts the performance of InterImage-XL close to the Swin-H-based SkySense (96.27 v.s. 96.32), which has more parameters. We also notice the accuracy of IMP-InterImage-XL is decreased marginally in EuroSAT after MTP. We will investigate this phenomenon later. Nevertheless, the obtained model still outperforms ${}^{+}$, which is pretrained with 1 million multispectral images, where each sample can be regarded as containing multiple groups of tri-spectral images, similar to RGB channels.
@@ IV-C Rotated Object Detection
IV-C1 Dataset
  1. 1)
  2. 2)
[…]
  3. 3)
  4. 4)
IV-C2 Implementation Details
IV-C3 Finetuning Results and Analyses
@@ IV-E3 Finetuning Results and Analyses
To comprehensively assess the finetuning performance of pretrained models, we conduct the comparison by collecting existing advanced change detection methods, as shown in Table VIII. It should be noted that, since the original WHU dataset does not provide an official train/test split, various split strategies are adopted in different methods. Therefore, on this dataset, we only list the accuracy value of the methods that employ the same settings as us or training with more images. It can be seen that MTP effectively improves the performances of pretrained models on these datasets. Especially, our models perform well on three large-scale datasets: WHU, LEVIR, and SVCD/CDD. Even if adopting simple UNet [76] and the RVSA model of the base version, the finetuning performances have been competitive and surpassed many advanced approaches. When utilizing larger models, the performance can be further boosted. Finally, they achieve the best accuracy on the WHU and LEVIR datasets by outperforming almost all existing methods, including the recent SkySense [42] that builds a larger change detection network with over 600M parameters, ChangeCLIP [184] that uses CLIP [17] to obtain additional knowledge from language modalities, and the newly proposed adapter BAN [185], where the ability of existing foundation model and change detection approaches can be exploited. Different from large-scale scenes, on the small-scale dataset OSCD, although MTP is still useful, the performances of our models have relatively large gaps compared to current works. These results suggest that it is necessary to conduct further explorations to enhance the model finetuning performance on datasets with small volumes and input sizes.
@@ IV-F3 Decoder Parameter Reusing
MTP utilizes task-specific decoders for segmentation and detection tasks. Hence, reusing these decoder weights during finetuning seems a naive choice. However, only semantic segmentation and rotated detection decoders are eligible for reuse, as per the segmentor or detector used in existing methods. We conduct experiments accordingly. Initially, during finetuning, aside from the backbone network, we initialize the corresponding decoders with pretrained weights. Employing ViT-B + RVSA, the results are presented in Table X. Across the six datasets, decoder parameter reusing (DPR) proves beneficial in only three scenarios. Notably, on segmentation tasks, DPR models fail to outperform the MTP models without DPR. Consequently, we conclude: after MTP, reusing pretrained decoder parameters in finetuning is unnecessary. Typically, decoders encode task-specific information. However, given that the SAMRS dataset used for pretraining involves annotations generated by SAM [47], they inevitably contain errors, jeopardizing the quality of pretrained decoders.

### v2

@@ II-C Multi-Task Pretraining for RS Foundation Model
Applying multi-task learning to enhance the RS foundation model is an intuitive idea. Li et al. [64] introduced multi-task SSL representation learning, combining image inpainting, transform prediction, and contrast learning to boost semantic segmentation performance in RS images. However, it was limited to finetuning a pretrained model solely on semantic segmentation tasks, constrained by model size and pretraining dataset capacity. RSCoTr [65] constructs a multi-task learning framework to simultaneously achieve classification, segmentation, and detection tasks. Unfortunately, the network can only be optimized by one task in each iteration during training due to lacking of multi-label datasets. The aspiration to consolidate multiple tasks into a single model has been a longstanding pursuit [15, 17, 66, 67, 68, 69, 70, 71, 72, 58, 73, 74, 42, 75], aligning with the original goals of the foundation model exploration. Bastani et al. [59] devised a multi-task model by integrating Swin-Base [57] with seven heads from existing networks (e.g., Faster-RCNN [76] and UNet [77]), facilitating training on the multi-task annotated Satlas dataset. However, their approach lacked incorporation of typical RS rotated object tasks, focusing solely on transferring the model to RS classification datasets. Inspired by these pioneering efforts, this paper constructs a unified multi-task pretraining framework that supports multiple datasets, where the sample in each dataset simultaneously possesses multi-task labels, which are uniformly processed in a data loader pipeline. During the training, the model can be simultaneously optimized through multiple tasks when in each iteration. Based on this framework, we pretrain RS foundation models with over 300M parameters, encompassing semantic segmentation, instance segmentation, and rotated object detection tasks using the SAMRS dataset. After pretraining, the backbone network is further finetuned on various RS downstream tasks.
@@ III-B Backbone Network
In this research, we adopt RVSA (Rotated Varied-Size window Attention) [30] and InternImage [82] as the representative vision transformer-based and CNN-based foundation models.
@@ III-B1 RVSA
This model is specially designed for RS images. Considering the various orientations of RS objects caused by the bird’s-eye view, this model extends the varied-size window attention in [83] by additionally introducing a learnable angle factor, offering windows that can adaptively zoom, translate, and rotate.
Specifically, given an input feature $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$ ($C,H,W$ are the number of channel, height, and width in $\mathbf{X}$), which is evenly divided into different windows, where the feature of each window can be formulated as $\mathbf{X}_{w}\in\mathbb{R}^{C\times s\times s}$ ($s$ is the window size), obtaining $\frac{H}{s}\times\frac{W}{s}$ windows totally. Then, three linear layers are used to generate the query feature, and initial key and value features, which are separately represented as $\mathbf{Q}_{w}$, $\mathbf{K}_{w}$ and $\mathbf{V}_{w}$. We use $\mathbf{X}_{w}$ to predict the variations of the window:
$S_{w},O_{w},\Theta_{w}=Linear(LeakyReLU(GAP(\mathbf{X}_{w})))$ (1)
Here, GAP is global average pooling, $S_{w}=\{s_{x},s_{y}\in\mathbb{R}^{1}\}$ and $O_{w}=\{o_{x},o_{y}\in\mathbb{R}^{1}\}$ are the scale factor and offset in the X and Y axis, while $\Theta_{w}=\{\theta\in\mathbb{R}^{1}\}$ is the rotation angle. Taking an example using the corner points of a window:
$\left[\begin{array}[]{c}x_{l/r}\\ y_{l/r}\\ \end{array}\right]=\left[\begin{array}[]{c}x^{c}\\ y^{c}\\ \end{array}\right]+\left[\begin{array}[]{c}x_{l/r}^{r}\\ y_{l/r}^{r}\\ \end{array}\right]$ (2)
where $x_{l},y_{l},x_{r},y_{r}$ are the coordinates of the upper left and lower right corners of the initial window, $x_{c},y_{c}$ are the coordinates of the window center point. Therefore, $x_{l}^{r},y_{l}^{r},x_{r}^{r},y_{r}^{r}$ are the distances between the corner points and the center in horizontal and vertical directions. The transformation of the window can be implemented using the obtained scaling, translation, and rotation factors:
$\begin{split}\left[\begin{array}[]{c}x_{l/r}^{\prime}\\ y_{l/r}^{\prime}\\ \end{array}\right]&=\left[\begin{array}[]{c}x^{c}\\ y^{c}\\ \end{array}\right]+\left[\begin{array}[]{c}o_{x}\\ o_{y}\\ \end{array}\right]\\ &+\left[\begin{array}[]{cc}\cos{\theta}&\sin{\theta}\\ -\sin{\theta}&\cos{\theta}\\ \end{array}\right]\left[\begin{array}[]{c}x_{l/r}^{r}\cdot s_{x}\\ y_{l/r}^{r}\cdot s_{y}\\ \end{array}\right]\end{split}$ (3)
$x_{l}^{\prime},y_{l}^{\prime},x_{r}^{\prime},y_{r}^{\prime}$ are the coordinates of the corner points of the transformed window. Then, new key and value feature $\mathbf{K}_{w^{\prime}}$ and $\mathbf{V}_{w^{\prime}}$ can be sampled from the obtained window, and the self-attention (SA) can be operated by the following formula:
$\mathbf{F}_{w}=SA(\mathbf{Q}_{w},\mathbf{K}_{w^{\prime}},\mathbf{V}_{w^{\prime}})=softmax(\frac{\mathbf{Q}_{w}\mathbf{K}_{w^{\prime}}^{T}}{\sqrt{C^{\prime}}})\mathbf{V}_{w^{\prime}}$ (4)
In this formula, $\mathbf{F}_{w}\in\mathbb{R}^{s^{2}\times C^{\prime}}$ is the feature output by one self-attention of one window, $C=hC^{\prime}$, where $h$ is the number of SA. The shape of final output features in RVSA can be recovered by concatenating the features from different SAs in the channel dimension and merging features from different windows along the spatial dimension.
RVSA is used to replace the original multi-head full attention in original vision transformers. To achieve a trade-off between accuracy and efficiency, following [84], only the full attention in 1/4 depth layer is preserved. In the original paper [30], RVSA is separately used on ViT [85] and ViTAE [72], whereas ViTAE is a CNN-Transformer hybrid model. In this paper, we employ the ViT-based version to investigate the impact of MTP on a plain vision transformer. In addition, the RVSA model in the original paper is limited to the base version of vision transformers, i.e., ViT-B + RVSA. To pretrain larger models, we further apply RVSA to ViT-Large, obtaining ViT-L + RVSA. Their detailed configurations are presented in Table I.
@@ III-B2 InternImage
III-C Multi-Task Pretraining
@@ III-D Implementation Details
IV Finetuning Experiments
In this section, we thoroughly evaluate MTP’s performance by finetuning pretrained models across four classical RS tasks: scene classification, object detection, semantic segmentation, and change detection, where the most representative and widely used benchmarks in the literature for different downstream tasks are employed for comparison. We also investigate the characteristics of MTP-based RS foundation models, examining the relationships between adopted datasets, hyperparameters, and finetuning performances, measuring accuracy variations with reduced training samples, and visualizing the predicted results.
@@ IV-A4 Finetuning Results and Analyses
Table IV shows the finetuning results. It can be seen that MTP can improve existing foundation models on scene classification tasks, especially for the RVSA series. It helps the model achieve state-of-the-art performances compared to other pretraining models that have comparable sizes. With the help of MTP, on the RESISC-45 dataset, InterImage-XL surpasses Swin-L-based SkySense [42], which is pretrained on a tremendously large dataset that has more than 20 million multimodal RS image triplets involving RGB high-resolution images and multi-temporal multispectral and SAR sequences. MTP boosts the performance of InterImage-XL close to the Swin-H-based SkySense (96.27 v.s. 96.32), which has more parameters. We also notice the accuracy of IMP-InterImage-XL is decreased marginally in EuroSAT after MTP. We will investigate this phenomenon later. Nevertheless, the obtained model still outperforms SpectralGPT+, which is pretrained with 1 million multispectral images, where each sample can be regarded as containing multiple groups of tri-spectral images, similar to RGB channels.
@@ IV-C Rotated Object Detection
IV-C1 Dataset
  1. 1)
  2. 2)
[…]
  3. 3)
  4. 4)
IV-C2 Implementation Details
IV-C3 Finetuning Results and Analyses
@@ IV-E3 Finetuning Results and Analyses
To comprehensively assess the finetuning performance of pretrained models, we conduct the comparison by collecting existing advanced change detection methods, as shown in Table VIII. It should be noted that, since the original WHU dataset does not provide an official train/test split, various split strategies are adopted in different methods. Therefore, on this dataset, we only list the accuracy value of the methods that employ the same settings as us or training with more images. It can be seen that MTP effectively improves the performances of pretrained models on these datasets. Especially, our models perform well on three large-scale datasets: WHU, LEVIR, and SVCD/CDD. Even if adopting simple UNet [77] and the RVSA model of the base version, the finetuning performances have been competitive and surpassed many advanced approaches. When utilizing larger models, the performance can be further boosted. Finally, they achieve the best accuracy on the WHU and LEVIR datasets by outperforming almost all existing methods, including the recent SkySense [42] that builds a larger change detection network with over 600M parameters, ChangeCLIP [188] that uses CLIP [17] to obtain additional knowledge from language modalities, and the newly proposed adapter BAN [189], where the ability of existing foundation model and change detection approaches can be exploited. Different from large-scale scenes, on the small-scale dataset OSCD, although MTP is still useful, the performances of our models have relatively large gaps compared to current works. We attribute the reason to data discrepancy and image size. Specifically, our models are pretrained on high-resolution RS images, which are similar to another three change detection datasets. However OSCD images are captured by multispectral sensors with lower resolutions. In addition, OSCD images are only cropped to 96 $\times$ 96 during training, which may be unsuitable for feature extraction, especially for non-hierarchical vision transformers. In the comparison methods, MATTER [24], SkySense [42], and GFM [46] use pyramid feature networks. Among them, MATTER and SkySense adopt Sentinel-2 multispectral image in pretraining, while GFM crops the OSCD image to a larger size, i.e., 192 $\times$ 192. In contrast, a relatively small image size (128 $\times$ 128) restricts the performances of ViT in SpectralGPT [34]. These results suggest that it is necessary to conduct further explorations to enhance the model finetuning performance on out-of-domain datasets with small volumes and input sizes.
@@ IV-F3 Decoder Parameter Reusing
MTP utilizes task-specific decoders for segmentation and detection tasks. Hence, reusing these decoder weights during finetuning seems a natural choice, and we conduct experiments accordingly using a backbone of ViT-B + RVSA. Specifically, aside from the backbone network, we also initialize the corresponding decoders with pretrained weights during finetuning. However, only semantic segmentation and rotated detection decoders are eligible for reuse, as per the segmentor or detector used in existing methods. Therefore, we performed the experiment on the corresponding six datasets, and the results have been presented in Table X. Among four detection datasets, decoder parameter reusing (DPR) proves beneficial in three scenarios, which are actually employed in performing MTP. Nevertheless, on another classical dataset DOTA-V1.0, which can be regarded as a subset of DOTA-V2.0, DPR is insufficient useful. On segmentation tasks, the performances of DPR models are decreased. We speculate that besides the domain differences compared to pretraining datasets, current models may be additionally affected by the segmentation labels in pretraining. This is because decoders typically encode task-specific information. However, given that the SAMRS dataset used for pretraining involves annotations generated by SAM [47], they inevitably contain errors, jeopardizing the quality of pretrained decoders. Finally, based on the above considerations, we conclude that after MTP, the performances of reusing pretrained decoder parameters in finetuning may depend on the similarity between pretraining and finetuning scenarios and the quality of pretraining annotations.